# A Test-Time Learning Approach to Reparameterize the Geophysical Inverse Problem with a Convolutional Neural Network

### v1

@@ II-D Necessary Explicit Regularization in DIP-Inv
One of the main purposes of this study is to test the implicit regularization effect introduced by the CNN structures. Although one might be inclined to try to eliminate all explicit regularization, this is a very ill-posed problem, and we find that including the smallness term is important for this problem. Thus, we choose to analyze the impacts of the implicit regularization effects of a CNN as compared to the role of the explicit smoothness terms in the conventional approach.
The other modification that we make is in how we treat the trade-off parameter $\beta$. The trade-off parameter $\beta$ in the proposed method is an exponential decay function.
@@
where t is the index for the epochs and $\tau$ is a constant decay rate. Namely, $\beta=e^{-\frac{i}{\tau}}$ in the i-th epoch. Although the conventional method uses a different cooling schedule and doesn’t employ $(1-\beta)$ in front of the $\phi_{d}$, we have found that this choice of cooling schedule performs well with the proposed method.
@@ II-E A practical implementation techniques
We choose Adam [19] as the optimizer for the DIP-Inv. Adam updates $w$ based on the gradient of the objective function, so we need to make sure that the gradient is correctly calculated. Let’s denote the objective function in 4 as $\mathcal{L}=(1-\beta)\phi_{d}+\beta\phi_{m}$. Note that we use the convention of column vector in this subsection.
@@
DIP-Inv uses SimPEG for the forward simulation and PyTorch for building the CNN architecture. We need to connect SimPEG (computing $J_{v}=\nabla_{m}\phi_{d}$) with PyTorch (automatic differentiate $w$). Here we play a trick: We define $\mathcal{L^{\prime}}=(1-\beta^{t})(J_{v}^{T}m^{t})+\beta^{t}||m^{t}-m_{ref}||^{1}$ and $J_{v}^{T}$ is considered as a fixed vector when calling .$\mathcal{L^{\prime}}.backward()$ and $optimizer.step()$ to update the weights $w$ in CNN. Although $L$ and $L^{\prime}$ are not identical, their gradients with respect to the weights $w$ are the same. Here, we elaborate on how we implement it:
@@
     using $simulation.Jtvec(m,W_{d}^{T}W_{d}(d^{pred}-d^{obs}))$
@@ III-A Survey Configuration and Mesh Design
We chose a dipole-dipole survey geometry with a station separation of 25 meters and a maximum of 24 receivers per transmitter. The survey line is 700 meters long, so the total number of potential difference measurements is 348 (= $3\times 24+(23+1)\times 23/2$).
For the simulation, we use a mesh that has core cells of 5m $\times$ 5m, with 200 core-mesh cells in the x-direction and 25 cells in the z-direction. We add 7 padding cells that extend by a factor of 1.5 to the sides and bottom of the mesh. In total, the mesh has $W\times H=6944$ cells. The total number of trainable weights in $L$ is 23055 based on the DIP-Inv architecture (see the Appendix for details), so we over-parameterize the problem. We also add a dropout layer to impose a stronger net regularization effect on this over-parameterized problem (see section IV-C). The subsurface conductivity models for Case 1 and Case 2 are shown in Fig. 3 and Fig. 4.
Figure 3: True Model for Case 1. The top layer has a conductivity of 0.02 S/m. The dike, with a dip angle of 45 degrees, has a conductivity of 0.1 S/m. The background has a conductivity of 0.01 S/m.
Figure 4: True Model for Case 2. The cylinder and the dike have a conductivity of 0.1 S/m. The background has a conductivity of 0.01 S/m.
@@ III-B Choice of optimization method
In the proposed method, we choose Adam, a first-order optimizer, with a learning rate of 0.0001 to update $w$ in $L$. Note that the conventional method uses Inexact Gaussian Newton [23], a second-order optimization method, to update $m$, since second-order optimization methods are known to speed up convergence for a variety of inverse problems[3]. However, researchers find practical challenges in employing second-order optimization methods to train a Neural Network (NN) partly because estimating the Hessian would be time-consuming in the case where the number of trainable parameters in an NN is large[8].
@@ III-C Case 1
We first find the best conventional inversion results with respect to the true model by conducting multiple conventional inversions using different values of $p_{s}$, $\alpha_{i}$, and $p_{i}$ (i = x or z) with or without sensitivity weighting.
@@
We find that the best conventional inversion results with respect to the true models are from $p_{s}=0$, $p_{x}=p_{z}=1$, $\alpha_{s}=0.005$, and $\alpha_{x}=\alpha_{z}=0.5$. See Fig. 5. We perform the proposed method with a decay rate of $\tau=1000$, and a dropout rate of 0.1 after the fourth hidden block.
@@
Figure 5: (a) is the true conductivity model for Case 1. (b) and (c) are the conventional inversion results without/with sensitivity weighting, respectively. (d) is the DIP-Inv result.
@@
All models fit the data to a comparable level, and the $\chi$-factor (the ratio of the final misfit with the target misfit, c.f. [22]) of each model is shown in the bottom right-hand-side of each inversion result. Though each shows a conductive, dipping structure, the details differ. In Figure 5b, we can see that by not including sensitivity weighting, we recover a more diffuse structure. We also see near-surface artifacts where we have high sensitivity to the electrode locations. Using sensitivity weighting, as shown in Figure 5c, reduces the near-surface artifacts and improves the recovery of the near-surface conductive layer. The dike structure is better defined, but the dip is more vertical than the true model. The DIP-Inv result (Figure 5d) does not require any sensitivity weighting and shows an improved recovery of the dip of the dike. We also recover the near-surface layer as a smooth structure.
To further test the recovery of the dip angle, we perform two other sets of inversions with two different true models that have different dip angles (See Case 1.2 and Case 1.3 in Fig. 6). The proposed method recovers the top layer well and surpasses the conventional method in terms of recovery of the dip angle.
@@
Figure 6: (a) is the true model; the dip of the dike is varied. (b) and (c) are from the conventional sparse-norms inversions without/with sensitivity weights respectively. (d) is the DIP-Inv result.
@@ III-D Case 2
We perform a set of conventional inversions and find that $p_{s}=0$, $p_{x}=p_{z}=2$, $\alpha_{s}=0.005$, and $\alpha_{x}=\alpha_{z}=0.5$ give the best conventional results with respect to the true model in Case 2. We then perform the proposed method with decay rate $\tau=800$, and the dropout rate 0.05 after the fourth hidden block. Again, all models fit the data to a comparable level. In Fig. 7b, c, the predicted models recover the cylindrical structure and a thin dike-like structure near the surface, but the structure of the dike doesn’t extend to depth. As a contrast, the DIP-Inv result (Fig. 7d) recovers the dike, and it is clear that it extends at depth, we also see the cylindrical target being recovered at the correct location, though its amplitude is reduced as compared to the true model.
Figure 7: (a) is the true model for Case 2.1. (b) and (c) are from the conventional sparse-norms inversions without/with sensitivity weights respectively. (d) is the DIP-Inv result.
To further verify the argument that the DIP-Inv method can distinguish the two compact anomalies in close proximity. We conduct another trial on a model with two compact, cylindrical targets (see Fig. 8a). Using the conventional approach without sensitivity weighting (Fig. 8b), we recover a blurred model that connects the two structures. Using sensitivity weighting is an improvement and we can see that there are two distinct anomalous bodies. They are somewhat smeared vertically (Fig. 8c). The DIP-Inv result arguably provides the best recovery of the two compact targets (Fig. 8d). Together, Cases 2.1 and 2.2 show that the DIP-Inv approach is beneficial for recovering compact targets that are in close proximity to one another.
Figure 8: (a) is the true model for Case 2.2. (b) and (c) are from the conventional sparse-norms inversions without/with sensitivity weights respectively. (d) is the DIP-Inv result.
@@ IV Implicit Regularization in DIP-Inv
In this section, we will explore the implicit regularization effect observed in the recovery models from the proposed method. We hypothesize that the smoothing effect generated by the bi-linear upsampling operator is a key component of the DIP-Inv approach. We test this hypothesis by replacing bi-linear upsampling with other upsampling operators. Finally, we present the effect of changing the number of hidden blocks and adding dropout layers.
@@ IV-A Bi-linear Upsampling Operator
In this subsection, we test the hypothesis that the observed implicit regularization effect is mainly from the bi-linear upsampling operator. The bi-linear upsampling is a common choice for up-scaling the dimension of certain hidden layers. If we consider that the $ith$ hidden layer is a bi-linear upsampling layer with upsampling scale $k$ and input dimension $(a,b,c,d)$, then the output dimension for the $ith$ hidden layer is $(a,b,kc,kd)$. The bi-linear interpolation is employed in this up-scaling process.
@@
The output value of the bi-linear interpolation in each pixel is a distance-based, weighted sum of the surrounding pixels; therefore, bi-linear interpolation generates a pixel-level smoothed output. To verify our hypothesis, we replace the bi-linear upsampling with the nearest neighbour upsampling or a transposed convolution. Our experiments demonstrated that these alternatives yielded inferior results (see Fig. 9 and 10), which supports that the bi-linear interpolation is critical in the success of DIP-Inv.
@@
Figure 9: (a) is the true model for Case 1.1. (b) and (c) are from inversion using nearest neighbour upsampling and transposed convolution operator respectively.
@@
Figure 10: (a) is the true model for Case 2.1. (b) and (c) are from inversion using nearest neighbour upsampling and transposed convolution operator respectively.
IV-B Choice of architecture
@@
To further test our hypothesis that the implicit regularization effect is partly from the bi-linear upsampling operator, we show the results from changing the number of bi-linear upsampling layers by two in the DIP-Inv. Since we argue that the implicit effect partially comes from the bi-linear upsampling layers, the regularization effect should be amplified as we increase the number of hidden blocks (one hidden block has one bi-linear upsampling layer), and vice versa. Similar to the need to choose a good trade-off between the data misfit term and the regularization term in the conventional method, we need to choose how much regularization effect we want to impose for our DIP-Inv method. Too much or too little regularization is both problematic.
@@
The DIP-Inv inversion results for Case 1 using 1, 3 and 5 hidden blocks are shown in Figure 11. We choose to not use a dropout layer in this subsection to avoid ambiguity on the source of the regularization effect. Using only one hidden block (Fig. 11b), we recover a model with lots of structure and artifacts, illustrating that one hidden block is insufficient for providing the regularization effect. On the other hand, when we use five hidden blocks, we recover a more blurred structure (Fig. 11d) because the additional hidden blocks provide more regularization. For this model, we find that using three hidden blocks (Fig.11c) provides an appropriate level of regularization for this model. Note that the optimal number of layers will depend upon the subsurface model. We repeat this experiment for the model from Case 2 to support our conclusion and show these results in Fig. 12. Again, we see that increasing the number of hidden blocks increases the regularization effect, and for this model, using three layers provides an appropriate level of regularization.
@@
Figure 11: (a) is the true model. (b), (c), and (d) are results for employing 1, 3, and 5 hidden blocks respectively. All results are from the CNN architectures without a dropout layer.
@@
Figure 12: (a) is the true model. (b), (c), and (d) are results for employing 1, 3, and 5 hidden blocks respectively. All results are from the CNN architectures without a Dropout layer.

### v2

@@ II-D Necessary Explicit Regularization in DIP-Inv
One of the main purposes of this study is to test the implicit regularization effect introduced by the CNN structures. Although one might be inclined to try to eliminate all explicit regularization, this is a very ill-posed problem (see Fig. 1 for example), and we find that including the smallness term is important for this problem. Thus, we choose to analyze the impacts of the implicit regularization effects of a CNN as compared to the role of the explicit smoothness terms in the conventional approach.
The other modification that we make is in how we treat the trade-off parameter $\beta$. The cooling schedule employed in the conventional method is not suitable for the DIP-Inv method due to the difference in the convergence rate of the two methods. Compared to the conventional method, the DIP-Inv method is searching for a solution in a higher dimension space, so the number of iterations needed to reach $\chi$-factor $\approx$ 1 is larger. As a result, we need a smoother and slower cooling schedule in the DIP-Inv method. The trade-off parameter $\beta$ in the DIP-Inv method is an exponential decay function.
@@
where t is the index for the epochs and $\tau$ is a constant decay rate. Namely, $\beta=e^{-\frac{i}{\tau}}$ in the i-th iteration.
@@ II-E A Practical Implementation Techniques
We choose Adam [39] as the optimizer for the DIP-Inv. Adam updates $w$ based on the gradient of the objective function, so we need to make sure that the gradient is correctly calculated. Let’s denote the objective function in 4 as $\mathcal{L}=(1-\beta)\phi_{d}+\beta\phi_{m}$. Note that we use the convention of column vector in this subsection. We make updates to $w$ by considering the gradient of $\mathcal{L}$ with respect to $w$. Here, we show the derivation for the i-th element in $w$.
@@
DIP-Inv uses SimPEG for the forward simulation and PyTorch for building the CNN architecture. We need to connect SimPEG, which computes $J_{v}=\nabla_{m}\phi_{d}$, with PyTorch which performs automatic differentiation with respect to $w$. To connect these two codes, we play a trick: We define $\mathcal{L^{\prime}}=(1-\beta^{t})(J_{v}^{T}m^{t})+\beta^{t}||m^{t}-m_{ref}||^{1}$ and $J_{v}^{T}$ is detached when calling .$\mathcal{L^{\prime}}.backward()$ and $optimizer.step()$ to update the weights $w$ in CNN. Although $\mathcal{L}$ and $\mathcal{L^{\prime}}$ are not identical, their gradients with respect to the weights $w$ are the same. Here, we elaborate on how we implement it:
@@
     using $simulation.Jtvec(m^{t},W_{d}^{T}W_{d}(d^{pred}-d^{obs}))$
@@ III-A Survey Configuration and Mesh Design
We choose a dipole-dipole survey geometry with a station separation of 25 meters and a maximum of 24 receivers per transmitter. The survey line is 700 meters long, so the total number of potential difference measurements is 348.
For the simulation, we use a mesh that has core cells of 5m $\times$ 5m, with 200 core-mesh cells in the x-direction and 25 cells in the z-direction. We add 7 padding cells that extend by a factor of 1.5 to the sides and bottom of the mesh. In total, the mesh has $W\times H=6944$ cells. The total number of trainable weights in $L$ is 23055 based on the DIP-Inv architecture [Table. I], so we over-parameterize the problem. We also add a dropout layer to impose a stronger net regularization effect on this over-parameterized problem (see section IV-C for further discussion). The subsurface conductivity models for Case 1 and Case 2 are shown in Fig. 4(a) and Fig. 5(a). To show how the current density changes due to the appearance of the anomalies, Fig. 4(b) and Fig. 5(b) illustrate the currents (streamlines) and resultant electric potentials (background colors) that are obtained by solving equation 1 for a single source with the positive electrode at location A and the negative electrode at location B. The positive electrode and negative electrode of the transmitter are labelled as A and B, respectively. The two electrodes of a receiver are labelled as M and N, respectively. We only demonstrate the location of one transmitter-receiver pair, the location of the other transmitter-receiver pairs can be referred from the survey configuration described above. The changes in current density are recorded in the voltage measurements on the ground. The voltage measurements are profiled on the pseudosections [Fig. 4(c) and Fig. 5(c)]. The pseudosection plots the apparent conductivity, calculated by a voltage measurement and the geometric factor which depends on the survey geometry, at the midpoint of the corresponding transmitter-receiver pair and at the depth that depends on the separation of the transmitter and receiver [3]. By measuring the potential difference between electrodes M and N shown in Fig. 4(b) or Fig. 5(b), we obtain a single datum that is shown as the white dot in the pseudosection. Although the pseudosection does not show a true geological crosssection, it is a useful way to visualize the measurements.
Figure 4: (a) is the true conductivity model for Case 1. The top layer has a conductivity of 0.02 S/m. The dike, with a dip angle of 45 degrees, has a conductivity of 0.1 S/m. The background has a conductivity of 0.01 S/m. (b) shows the electrical potential and current density with the 6th transmitter for Case 1. (c) shows the pseudosection for Case 1.
Figure 5: (a) is the true conductivity model for Case 2. The cylinder and the dike have a conductivity of 0.1 S/m. The background has a conductivity of 0.01 S/m. (b) shows the electrical potential and current density with the 14th transmitter for Case 2. (c) shows the pseudosection for Case 2.
@@ III-B Choice of Optimization Method
In the DIP-Inv method, we choose Adam, a first-order optimizer, with a learning rate of 0.0001 to update weights $w$ in $L$. Note that we only have one input ”sample”, which is a non-trainable 1 by 8 tensor, so the objective function is not stochastic in this study. The conventional method uses inexact Gauss Newton [40], a second-order optimization method, to update $m$ since the second-order optimization methods are known to speed up convergence for a variety of inverse problems [5]. However, researchers found practical challenges in employing second-order optimization methods to train a NN partly because estimating the Hessian is time-consuming when the number of trainable parameters is large [41]. As a result, we employ inexact Gauss Newton to update $m$, whose dimension is 6944, in the conventional method and Adam to update $w$, whose dimension is 23055, in the DIP-Inv method.
@@ III-C Case 1
We first find the best conventional inversion results with respect to the true model by a grid search over the hyper-parameter space. Namely, we conduct multiple conventional inversions using different values of $p_{s,x,z}$, $\beta_{s}$, $\alpha_{s,x,z}$ with or without sensitivity weighting.
@@
After that, we find that the best conventional inversion results with respect to the true model are from $\beta_{s}=1e2$, $p_{s}=0$, $p_{x}=p_{z}=1$, $\alpha_{s}=0.005$, and $\alpha_{x}=\alpha_{z}=0.5$ [Fig. 6(b) and (c)]. We perform the DIP-Inv method with a decay rate of $\tau=1000$, and a dropout rate of 0.1 after the fourth hidden block [Fig. 6(d)].
@@
Figure 6: (a) is the true conductivity model. (b) and (c) are the conventional inversion results without/with sensitivity weighting, respectively. (d) is the DIP-Inv result.
@@
All models fit the data to a comparable level, and the $\chi$-factor of each model is shown in the bottom right-hand side of each inversion result. Though each shows a conductive, dipping structure, the details differ. By not including sensitivity weighting, we recover a more diffuse structure [Fig. 6(b)]. We also see near-electrodes artifacts. Using sensitivity weighting, as shown in Fig. 6(c), reduces the near-electrodes artifacts, but the dip is more vertical than the true model. The DIP-Inv result [Fig. 6(d)] does not require any sensitivity weighting and shows an improved recovery of the dip of the dike without any near-electrodes artifacts. In a nutshell, the predicted model in Fig. 6(b) has a better indication of dip but has near-surface artifacts, whereas Fig. 6(c) removes those artifacts but has the wrong dip angle; therefore, each has useful information, but neither of them satisfactory on its own. In contrast, the predicted model in Fig. 6(d) has good dip angle prediction and no near-electrodes artifacts.
To further test the recovery of the dip angle, we perform two other sets of inversions with two different true models that have different dip angles (See Case 1.2 and Case 1.3 in Fig. 7). In the conventional inversion results without using sensitivity weighting, anomaly predictions on the top layer concentrate near the electrodes, and the conductivity values predicted in the dike locations are lower than the true values. The conventional inversion results using sensitivity weighting predict unwanted diffuse structures on the bottom of the dike, especially for Case 1.2. In comparison, the DIP-Inv method recovers the top layer well and surpasses the conventional method in terms of recovery of the dip angle. The conductivity values along a depth profile in Fig. 8 show that the inversion result from the DIP-Inv method follows a more similar trend as the true model compared to the inversion result from the conventional method in all 3 cases: The DIP-Inv method improves the recovery of the position of the dike as compared to the conventional method where it is recovered in a deeper position than the true model. We also show the mean absolute error (MAE) and the mean squared error (MSE) between the inversion results and the true models [Table. II], which are evaluation metrics that are common in the ML community to compare the performance of algorithms. For the conventional method, we show the scores for the models with sensitivity weighting. The DIP-Inv method has better performance in both metrics on all three models (Case 1.1, 1.2, and 1.3). Considering that the inversion results from the conventional method and the DIP-Inv method have similar chi-factor values, we conclude that the DIP-Inv method produces better recovery of the subsurface structures than the conventional method when both inversions fit observations at the same level.
@@
Figure 7: (a) is the true model; the dip of the dike is varied. (b) and (c) are from the conventional sparse-norms inversions without/with sensitivity weighting respectively. (d) is the DIP-Inv result.
Figure 8: The log conductivity values along the vertical profile at x = -180m for Case 1.1-1.3. The locations of these profiles are marked by the white vertical lines in Fig. 7(a). Only conventional results from the inversions with sensitivity weighting are shown in the plot.
TABLE II: Comparison of the DIP-Inv method and the conventional method with sensitivity weighting
Model Metrics Conventional method DIP-Inv
Case 1.1 MAE $\downarrow$ 0.1862 0.1500
MSE $\downarrow$ 0.0058 0.0051
Case 1.2 MAE $\downarrow$ 0.1832 0.1524
MSE $\downarrow$ 0.0057 0.0051
Case 1.3 MAE $\downarrow$ 0.1691 0.1564
MSE $\downarrow$ 0.0059 0.0054
@@ III-D Case 2
After a grid search over the hyper-parameter space, we find that $\beta_{s}=1e1$, $p_{s}=0$, $p_{x}=p_{z}=2$, $\alpha_{s}=0.005$, and $\alpha_{x}=\alpha_{z}=0.5$ give the best conventional results with respect to the true model in Case 2 [Fig. 9(b), (c)]. We then perform the DIP-Inv method with decay rate $\tau=800$, and a dropout rate of 0.05 after the fourth hidden block [Fig. 9(d)].
Again, all models fit the data to a comparable level. In Fig. 9(b) and (c), the predicted models recover the cylindrical structure and a thin dike-like structure near the surface, but the structure of the dike doesn’t extend to depth. In contrast, the DIP-Inv result [Fig. 9(d)] recovers the dike, and it is clear that it extends at depth, we also see the cylindrical target being recovered at the correct location, though its amplitude is reduced as compared to the true model.
Figure 9: (a) is the true model. (b) and (c) are from the conventional sparse-norms inversions without/with sensitivity weighting respectively. (d) is the DIP-Inv result.
To further verify the argument that the DIP-Inv method can distinguish the two compact anomalies in close proximity. We conduct another trial on a model with two compact, cylindrical targets [Fig. 10a]. Using the conventional approach without sensitivity weighting [Fig. 10(b)], we recover a blurred model that connects the two structures. Using sensitivity weighting is an improvement and we can see that there are two distinct bodies. They are somewhat smeared vertically [Fig. 10(c)]. The DIP-Inv result arguably provides the best recovery of the two compact targets [Fig. 10(d)].
[…]
Figure 10: (a) is the true model. (b) and (c) are from the conventional sparse-norms inversions without/with sensitivity weighting respectively. (d) is the DIP-Inv result.
TABLE III: Comparison of the DIP-Inv method and the conventional method with sensitivity weighting
Model Metrics Conventional method DIP-Inv
Case 2.1 MAE $\downarrow$ 0.1313 0.0997
MSE $\downarrow$ 0.0043 0.0047
Case 2.2 MAE $\downarrow$ 0.0533 0.0580
MSE $\downarrow$ 0.0039 0.0033
@@ IV Implicit Regularization in DIP-Inv
In this section, we will explore the implicit regularization effect observed in the results from the DIP-Inv method. We hypothesize that the smoothing effect generated by the bi-linear upsampling operator is a key component of the DIP-Inv approach. We then test this hypothesis by replacing bi-linear upsampling with other upsampling operators. Finally, we present the effect of changing the number of hidden blocks and adding dropout layers.
@@ IV-A Bi-linear Upsampling Operator
In this subsection, we test the hypothesis that the observed implicit regularization effect is mainly from the bi-linear upsampling operator. Bi-linear upsampling is a common choice for up-scaling the dimension of certain hidden layers. If we consider that the i-th hidden layer is a bi-linear upsampling layer with upsampling scale $k$ and input dimension $(a,b,c,d)$, then the output dimension for the i-th hidden layer is $(a,b,kc,kd)$. The bi-linear interpolation is employed in this up-scaling process.
@@
The output value of the bi-linear interpolation in each pixel is a distance-based, weighted sum of the surrounding pixels; therefore, bi-linear interpolation generates a pixel-level smoothed output. To verify our hypothesis, we replace the bi-linear upsampling with the nearest neighbour upsampling or the transposed convolution. Our experiments demonstrated that these alternatives yield inferior results [Fig. 11 and 12], which supports that the bi-linear upsampling is critical in the success of DIP-Inv.
@@
Figure 11: (a) is the true model. (b) and (c) are from inversions using nearest neighbour upsampling and transposed convolution operator respectively.
@@
Figure 12: (a) is the true model. (b) and (c) are from inversions using nearest neighbour upsampling and transposed convolution operator respectively.
IV-B Influence of Number of Layers
@@
To further test our hypothesis that the implicit regularization effect is partly from the bi-linear upsampling operator, we show the results from changing the number of bi-linear upsampling layers by two. If our hypothesis is true, then the regularization effect should be amplified as we increase the number of hidden blocks (one hidden block has one bi-linear upsampling layer), and vice versa. Similar to the need to choose a good trade-off between the data misfit term and the regularization term in the conventional method, we need to choose how much regularization effect we want to impose for our DIP-Inv method on a case-by-case basis. Too much or too little regularization is both problematic.
@@
The DIP-Inv inversion results for Case 1 using 1, 3 and 5 hidden blocks are shown in Fig. 13. We don’t use a dropout layer for all experiments presented in this subsection to avoid ambiguity on the source of the regularization effect. Using only one hidden block [Fig. 13(b)], we recover a model with lots of structure and artifacts, illustrating that one hidden block is insufficient for providing enough regularization effect. On the other hand, when we use five hidden blocks, we recover a more blurred structure [Fig. 13(d)] because the additional hidden blocks provide too much smooth effect. For this model, we find that using three hidden blocks [Fig.13(c)] provides an appropriate level of regularization. Note that the optimal number of layers will depend on the subsurface model. We repeat this experiment for the model from Case 2 to support our conclusion and show these results in Fig. 14. Again, we see that increasing the number of hidden blocks increases the regularization effect and using three layers provides an appropriate level of regularization.
@@
Figure 13: (a) is the true model. (b), (c), and (d) are results for employing 1, 3, and 5 hidden blocks respectively. All results are from the CNN architectures without any dropout layer.
@@
Figure 14: (a) is the true model. (b), (c), and (d) are results for employing 1, 3, and 5 hidden blocks respectively. All results are from the CNN architectures without any dropout layer.